Geraldine uses a 439-nm coherent light source and a double slit with a slit separation of 26.0 μm to display three interference maxima on a screen that is 20.0 cm wide. If she wants to spread the three bright fringes across the full width of the screen, from a minimum on one side to a minimum on the other side, how far from the screen should she place the double slit?
1
Expert's answer
2020-05-04T12:44:18-0400

As per the given question,

wavelength of the wave "(\lambda)=439nm"

separation between the slit "(d)=26\mu m"

width of the interference maxima =20.0 cm

Hence "20\/2=10cm"

total number of interference maxima (m)=3

Let the initial distance between the slit and the screen is D

now, we know that "y=\dfrac{m\lambda D}{d}"

"D=\dfrac{yd}{m\lambda}=\dfrac{10\times 10^{-2}\times 26\times 10^{-6}}{3\times 439\times10^{-9}}=1.97m \simeq 2m"

Now, for the second case y=20,

m=3

"(d)=26\mu m"

"D=\dfrac{yd}{m\lambda}=\dfrac{20\times 10^{-2}\times 26\times 10^{-6}}{3\times 439\times10^{-9}}=3.95m \simeq 4m"